OpenStudy (anonymous):

(Help Please :D )Mr. Chavez spent $199.99 for a new smart phone, $39.99 on a carrying case, and $59.99 on accessories. The expression |-199.00| + |-39.99| + |-59.99| represents the total amount that Mr. Chavez spent. How much did Mr. Chavez spend altogether? Check your answer using estimation. (Is the answer 299.97?)

OpenStudy (anonymous):

yes that is correct if you didn't know since 199.99 + 99.98 = 299.97 can I get a medal?

OpenStudy (anonymous):

$59.99 + 39.99 = 99.98 +199.99 = 299.97

OpenStudy (anonymous):

@mathmale is this correct? (just making sure...)

OpenStudy (anonymous):

@TheSmartOne @Mehek14